E. 90 31° 23. A rescue airplane is diving at an angle of 370 below the horizontal with a speed of 250 m/s. It releases a survival package when it is at an altitude of 600 m. If air resistance is ignored, the horizontal distance of the point of impact from the plane at the moment of the packages release is A. 2.80团103 m. B. 720 m. C. 6.80 2 103 m. D. 420 m. E. 5.50 103 m.
Answer #1

here,

theta = 37 degree

initial speed of plane , u = 250 m/s

let the time taken be t

for vertical direction

h = u * sin(theta) * t + 0.5 * g * t^2

600 = 250 * sin(37) * t + 0.5 * 9.81 * t^2

solving for t

t = 3.57 s

the horizontal distance travelled , x = u * cos(theta) * t

x = 250 * cos(37) * 3.57 m

x = 720 m

the horizontal distance travelled is B) 720 m